OpenStudy (anonymous):

By selling 50 pens for Rs.160 a person loses 20%. The number of pens he should sell for Rs. 96 so as to have a profit of 20% is a) 20 b) 16 c) 24 d) none of the above

OpenStudy (anonymous):

sells each pen for 160/50 = 3.2 Rs. he loses 20% so cp was 0.32 / 0.8 = 40 R to make 20 prpfit he should sell ay 48 Rs for 96 he sells 96 / 4.8 = 20 answer a)